# A hybrid electromechanical phase-field and deep learning framework for predicting fracture in dielectric nanocomposites

### v1

@@ Abstract
The accurate and efficient prediction of crack propagation in dielectric materials is a critical challenge in structural health monitoring and the design of smart systems. This work presents a hybrid modeling framework that combines an electromechanical phase-field fracture model with deep learning-based surrogate modeling to predict fracture evolution in dielectric nanocomposite plates. The underlying finite element simulations capture the coupling between mechanical deformation and electrical field perturbations caused by cracks, using a variational phase-field formulation. High-fidelity simulation outputs–namely, phase-field damage variables and electric potential fields–are used to train convolutional neural networks (CNNs) with ResNet-U-Net architectures for pixel-wise segmentation of crack paths. The study systematically compares the performance of CNNs trained on phase-field versus electric potential data across multiple ResNet backbones. The results reveal that electric potential fields, although they encode damage indirectly, offer superior segmentation accuracy, faster convergence, and enhanced generalization, owing to their smoother gradient distribution and global spatial coverage. The proposed framework significantly reduces computational costs while preserving high accuracy, offers potential when appropriately adapted for sensor-based input data.
@@
A. Electroactive fracture, B. Phase-field modeling, C. Electrical potential, D. Convolutional neural networks, E. Dielectric nanocomposites, F. Structural health monitoring
@@ 1 Introduction
The relentless pursuit of higher performance, efficiency, and safety in modern engineering has led to the widespread adoption of advanced materials, among which polymer composites stand preeminent. Their integration into critical sectors, including the aerospace, automotive, naval, and biomedical industries, is driven by an exceptional combination of high specific stiffness and strength, low density, and superior fatigue endurance when compared to traditional monolithic materials like metal alloys, see [1, 2, 3]. The demand for these materials continues to grow at an exponential rate, a trend underscored by their expanding utilization in primary aircraft structures and other high-stakes industrial applications [4]. However, the full potential of composites is often tempered by the profound complexity of their failure mechanisms, a direct consequence of their inherent heterogeneity and anisotropy [5].
@@
The paradigm of composite materials is currently undergoing another significant evolution with the advent of multifunctional or ”smart” composites, which are engineered to perform tasks beyond their primary structural role, such as sensing, actuation, or self-healing, see [6, 7, 8] among others. A particularly promising class of these materials for structural health monitoring (SHM) applications is the dielectric nanocomposite. These are typically formed by embedding dielectric nanofillers, such as copper phthalocyanine (CuPc), into a polymer matrix to create a dielectric elastomer–a type of electroactive polymer that exhibits strong electromechanical coupling [9]. This coupling manifests in two ways: the material deforms when subjected to an electric field, and conversely, its electrical properties–notably its dielectric permittivity and capacitance–change in response to mechanical strain or damage [10]. This latter effect forms the basis of their application in SHM. By continuously monitoring the electrical response of the material, it becomes possible to detect, locate, and even quantify the extent of damage in real-time, creating a ”self-sensing” structure that can diagnose its own health condition [8, 11]. This capability obviates the need for many traditional, often cumbersome and offline, non-destructive testing (NDT) techniques and is paramount for ensuring the operational safety and reliability of structures where failure could be catastrophic. This integration of sensing functionality directly into the load-bearing material introduces a fundamental duality that poses a profound modeling challenge. In these dielectric nanocomposites, the material is simultaneously the structure and the sensor [11]. A fracture event, such as the initiation and propagation of a crack, is therefore not merely a mechanical failure; it is intrinsically a coupled electromechanical phenomenon. The crack degrades the mechanical integrity of the composite by reducing its stiffness and strength, while at the same time, it perturbs the local electrical field by altering the dielectric constant in the damaged region, thereby directly impacting the sensing functionality, see [12, 13]. Consequently, any predictive model that aims for physical accuracy cannot treat these aspects in isolation. A simple mechanical fracture model is insufficient because it neglects the electrical response, and a simple electrical model is insufficient because it ignores the mechanical failure driving the change. A fully coupled electromechanical fracture model is therefore not just an academic exercise but a fundamental necessity for understanding the behavior, predicting the reliability, and guiding the design of these advanced multifunctional material systems.
@@
Given the intricate nature of damage evolution in composites, purely experimental characterization is often insufficient for achieving reliable design and life prediction [14]. This has established computational modeling as an indispensable tool in the field of fracture mechanics. Over the past two decades, the phase-field (PF) method has emerged as a particularly powerful and versatile computational framework, overcoming many of the significant algorithmic and implementation challenges associated with classical discrete fracture approaches, such as the extended finite element method (XFEM), see [15, 16, 17, 18, 19, 20, 21, 22] among others. The core concept of PF method is to abandon the explicit representation of cracks as sharp geometric discontinuities. Instead, a crack is represented in a diffuse or ”smeared” manner over a narrow region by introducing a continuous scalar variable, known as the phase field, typically denoted by $\mathfrak{d}$’. This field acts as an order parameter, smoothly transitioning from an undamaged state (e.g., $\mathfrak{d}=0$’) to a fully fractured state ($\mathfrak{d}=1$’), see [23, 24, 25] and the references cited therein. This diffusive representation is governed by a partial differential equation derived from a variational energy principle. This allows the model to naturally and robustly predict complex fracture phenomena, including the initiation of cracks from arbitrary sites, their subsequent propagation along complex paths, and events like crack branching and coalescence, all without the need for ad-hoc criteria or complex topological tracking and remeshing algorithms.
@@
The theoretical underpinnings of PF method are robust, originating from the variational theory of brittle fracture proposed by Francfort and Marigo, which recasts fracture as a global energy minimization problem [23]. The phase-field formulation can be rigorously shown to be a regularized approximation of this foundational theory [24]. The great flexibility of the PF method framework, where the system’s behavior is governed by an energy functional, allows for its extension to coupled multi-physics problems by incorporating additional energy terms, see [13, 16, 26]. This capability has been leveraged to develop phase-field models for fracture in piezoelectric and ferroelectric materials, which, like the dielectric nanocomposites of interest here, exhibit strong electromechanical coupling [12, 13]. These advanced models couple the variational formulation of fracture with the governing equations of electrostatics. A key challenge in this domain is the accurate modeling of the electrical boundary conditions on the crack faces, as the behavior depends on the electrical properties of the medium (e.g., air, vacuum) filling the crack gap. Different physical scenarios must be considered, such as an electrically impermeable (insulating) crack, a permeable crack, or a semi-permeable crack [27]. In the PF method framework, these conditions are encoded by selectively degrading different terms within the total electromechanical enthalpy density function based on the value of the phase-field variable [12, 13]. A physically consistent model must therefore capture the simultaneous degradation of both the mechanical stiffness and the dielectric permittivity as the crack propagates, introducing significant nonlinearity and mathematical complexity into the governing equations.
@@
While the phase-field method provides the necessary physical fidelity to capture these complex, coupled phenomena, this power comes at a steep and often prohibitive price: extreme computational cost [28]. This ”curse of fidelity” arises from several compounding factors. First, a standard PF method simulation already necessitates the use of a very fine computational mesh to accurately resolve the narrow, diffuse crack region, as the mesh size must be smaller than the phase-field length scale parameter. Second, when coupling with another physical field like electrostatics, a new set of governing equations and variables (e.g., electric potential) must be solved at every node in the mesh, increasing the size of the algebraic system. Third, and most critically, the coupling terms in the energy functional introduce strong nonlinearities, requiring more iterations within each time step for the numerical solver to converge. The combination of these factors leads to a multiplicative, rather than additive, increase in computational expense. As a result, high-fidelity, coupled electromechanical phase-field simulations can be so computationally intensive that they become impractical for tasks such as large-scale structural analysis, iterative design optimization, or real-time damage assessment, which are precisely the applications for which smart materials are intended. This establishes a critical research gap: a powerful theoretical tool exists to model the physics correctly, but it is too slow to be of practical use. There is a pressing need for a new computational paradigm that can retain the physical accuracy of coupled PF method while drastically reducing the associated computational burden.
@@
To surmount the computational bottleneck imposed by high-fidelity simulations, the scientific community is increasingly turning to machine learning (ML) and deep learning (DL). The central strategy is to develop a data-driven ”surrogate” model–a trained neural network that learns to approximate the complex input-output relationship of the expensive physics-based simulation but can make predictions at a fraction of the computational cost [29, 30, 31, 32, 33]. The application of ML, particularly artificial neural networks (ANNs), to problems in fracture mechanics is a rapidly expanding field, see [29, 30, 34, 35] among others. Early efforts have demonstrated the ability of ANNs to predict key fracture parameters, such as the stress intensity factor (SIF) for cracked plates, by training the networks on large datasets generated via conventional finite element analysis (FEA).
@@
For problems in continuum mechanics, where the outputs are spatially distributed fields (e.g., stress, strain, displacement, or phase-field), convolutional neural networks (CNNs) have emerged as a particularly suitable architecture. Originally developed for and famously successful in the domain of image recognition, CNNs are inherently designed to process grid-like data and learn hierarchical spatial features, see [36, 37, 38] and the references cited therein. The convolutional layers of a CNN act as learnable filters that slide across the input data, identifying local patterns. This mechanism is conceptually analogous to how physical phenomena, such as stress concentrations, manifest locally within a material domain, for instance, around the tip of a crack [39]. The efficacy of this approach has been demonstrated in several studies where CNNs have been successfully employed as surrogate models to predict the effective mechanical properties of periodic composites, see [40, 41, 42] among many others. In these works, the material microstructure is represented as a binary image, and the CNN is trained to regress the homogenized properties, achieving prediction speeds that are orders of magnitude faster than direct FEA simulations.
@@
The accurate and efficient prediction of fracture in multifunctional dielectric nanocomposites is essential for their reliable design and deployment in next-generation technologies, particularly for real-time structural health monitoring. While coupled electromechanical phase-field models possess the requisite physical fidelity to capture the complex failure behavior of these materials, their prohibitive computational expense renders them impractical for the very applications they are meant to enable. This creates a critical disconnect between modeling capability and practical utility. To bridge this gap, this study proposes and validates a novel hybrid computational framework that synergistically combines high-fidelity, fully coupled electromechanical phase-field fracture modeling with deep learning-based surrogate modeling. By training a deep neural network on data generated from the physics-based PF method, the framework aims to achieve the predictive accuracy of the high-fidelity model at a computational speed amenable to large-scale analysis and design. The primary objectives and specific contributions of this paper are fourfold: (i) to develop and implement a thermodynamically consistent electromechanical phase-field fracture model for dielectric nanocomposites, capable of capturing the coupled degradation of mechanical stiffness and dielectric permittivity under electromechanical loading, (ii) to utilize this high-fidelity simulation framework to generate a comprehensive, high-resolution dataset of simulation results. This data set, which encompasses the evolution of phase-field and electric potential variables, serves as the ground-truth data for training the deep learning surrogate model, (iii) to design and train a deep convolutional neural network (CNN) based on ResNet-U-Net architectures to learn the evolution of the coupled physical fields, and (iv) to conduct a rigorous evaluation of the surrogate model’s performance, assessing its predictive accuracy on unseen scenarios and generalization capability.

### v2

@@ Abstract
The accurate and efficient prediction of crack propagation in dielectric materials is a critical challenge in structural health monitoring and the design of smart systems. This work presents a hybrid modeling framework that combines an electromechanical phase-field fracture model with deep learning-based surrogate modeling to predict fracture evolution in dielectric nanocomposite plates. The underlying finite element simulations capture the coupling between mechanical deformation and electrical field perturbations caused by cracks, using a variational phase-field formulation. High-fidelity simulation outputs–namely, phase-field damage variables and electric potential fields–are utilized to train convolutional neural networks (CNNs) with ResNet-U-Net architectures. Crucially, the framework is designed to predict the final crack path directly from the geometric configuration (random defect patterns). We systematically compare the effectiveness of using either phase-field variables or electric potential fields as the primary physical signatures to guide the training process. The results reveal that models informed by electric potential fields offer superior segmentation accuracy, faster convergence, and enhanced generalization, owing to the smoother gradient distribution and global spatial coverage of the electrical response. Ultimately, the trained surrogate model enables the instantaneous, geometry-driven prediction of crack paths, bypassing the need for computationally intensive field calculations during inference. This demonstrates that leveraging electrical physics as a ”training guide” significantly improves the reliability of real-time fracture assessment in smart materials.
@@
A. Electroactive fracture, B. Phase-field modeling, C. Electric potential, D. Convolutional neural networks, E. Dielectric nanocomposites, F. Structural health monitoring
@@
Nomenclature
Symbols
$\mathbf{u}$ Displacement vector $\boldsymbol{\varepsilon}$ Infinitesimal strain tensor $\boldsymbol{\sigma}$ Cauchy stress tensor $\mathbf{f}$ Body force vector $\phi$ Electric potential $\mathbf{E}$ Electric field vector $\mathbf{D}$ Electric displacement vector $\rho_{e}$ Free charge density $\mathbf{t}$ Mechanical traction $q_{e}$ Electric surface charge density $\mathfrak{d}$ Phase-field damage variable $\ell$ Regularization length scale parameter $G_{c}$ Critical energy release rate / fracture toughness $\psi^{\text{mech}}$ Elastic energy density $\psi^{\text{elec}}$ Electrostatic energy density $\psi^{\text{frac}}$ Fracture energy density $\mathbb{C}$ Fourth-order elasticity tensor $\lambda,\mu$ Lamé parameters $E$ Young’s modulus $\nu$ Poisson’s ratio $\boldsymbol{\epsilon}$ Dielectric permittivity tensor $\epsilon$ Dielectric permittivity $\mathcal{H}$ History field for irreversibility
Functions and parameters
$g_{1}$ Mechanical degradation function $g_{2}$ Dielectric degradation function $\kappa_{\eta}$ Residual stiffness parameter $\kappa_{\epsilon}$ Residual permittivity parameter $k,n$ Degradation parameters
Abbreviations
CNN Convolutional neural network U-Net Encoder–decoder convolutional neural network ResNet Residual neural network FE Finite element SHM Structural health monitoring
@@ 1 Introduction
The relentless pursuit of higher performance, efficiency, and safety in modern engineering has led to the widespread adoption of advanced materials, among which polymer composites stand preeminent. Their integration into critical sectors, including the aerospace, automotive, naval, and biomedical industries, is driven by an exceptional combination of high specific stiffness and strength, low density, and superior fatigue endurance when compared to traditional monolithic materials like metal alloys, see Irving and Soutis [2020], Zhang et al. [2018], Bhong et al. [2023]. The demand for these materials continues to grow at an exponential rate, a trend underscored by their expanding utilization in primary aircraft structures and other high-stakes industrial applications Stoll and Weidenmann [2018]. However, the full potential of composites is often tempered by the profound complexity of their failure mechanisms, a direct consequence of their inherent heterogeneity and anisotropy Zhang et al. [2023].
@@
The paradigm of composite materials is currently undergoing another significant evolution with the advent of multifunctional or ”smart” composites, which are engineered to perform tasks beyond their primary structural role, such as sensing, actuation, or self-healing, see Kontiza and Kartsonakis [2024], Lee et al. [2024], Thorn et al. [2023] among others. A particularly promising class of these materials for structural health monitoring (SHM) applications is the dielectric nanocomposite. These are typically formed by embedding dielectric nanofillers, such as copper phthalocyanine (CuPc), into a polymer matrix to create a dielectric elastomer–a type of electroactive polymer that exhibits strong electromechanical coupling Hu et al. [1549]. This coupling manifests in two ways: the material deforms when subjected to an electric field, and conversely, its electrical properties–notably its dielectric permittivity and capacitance–change in response to mechanical strain or damage Pelrine et al. [2000]. This latter effect forms the basis of their application in SHM. By continuously monitoring the electrical response of the material, it becomes possible to detect, locate, and even quantify the extent of damage in real-time, creating a ”self-sensing” structure that can diagnose its own health condition Thorn et al. [2023], Hassan and Tallman [2022]. This capability obviates the need for many traditional, often cumbersome and offline, non-destructive testing (NDT) techniques and is paramount for ensuring the operational safety and reliability of structures where failure could be catastrophic. This integration of sensing functionality directly into the load-bearing material introduces a fundamental duality that poses a profound modeling challenge. In these dielectric nanocomposites, the material is simultaneously the structure and the sensor Hassan and Tallman [2022]. A fracture event, such as the initiation and propagation of a crack, is therefore not merely a mechanical failure; it is intrinsically a coupled electromechanical phenomenon. The crack degrades the mechanical integrity of the composite by reducing its stiffness and strength, while at the same time, it perturbs the local electrical field by altering the dielectric constant in the damaged region, thereby directly impacting the sensing functionality, see Zhang et al. [2022], Quinteros et al. [2023]. Consequently, any predictive model that aims for physical accuracy cannot treat these aspects in isolation. A simple mechanical fracture model is insufficient because it neglects the electrical response, and a simple electrical model is insufficient because it ignores the mechanical failure driving the change. A fully coupled electromechanical fracture model is therefore not just an academic exercise but a fundamental necessity for understanding the behavior, predicting the reliability, and guiding the design of these advanced multifunctional material systems.
@@
Given the intricate nature of damage evolution in composites, purely experimental characterization is often insufficient for achieving reliable design and life prediction Tan et al. [2015]. This has established computational modeling as an indispensable tool in the field of fracture mechanics. Over the past two decades, the phase-field (PF) method has emerged as a particularly powerful and versatile computational framework, overcoming many of the significant algorithmic and implementation challenges associated with classical discrete fracture approaches, such as the extended finite element method (XFEM), see Kristensen et al. [2021], Zhuang et al. [2022], Carollo et al. [2018], Dean et al. [2020b, a], Asur Vijaya Kumar et al. [2021a, c], Bui and Hu [2021], Kumar et al. [2025], Asur Vijaya Kumar et al. [2021b] among others. The core concept of PF method is to abandon the explicit representation of cracks as sharp geometric discontinuities. Instead, a crack is represented in a diffuse or ”smeared” manner over a narrow region by introducing a continuous scalar variable, known as the phase field, typically denoted by $\mathfrak{d}$’. This field acts as an order parameter, smoothly transitioning from an undamaged state (e.g., $\mathfrak{d}=0$’) to a fully fractured state ($\mathfrak{d}=1$’), see Francfort and Marigo [1998], Bourdin et al. [2000], Miehe et al. [2015] and the references cited therein. This diffusive representation is governed by a partial differential equation derived from a variational energy principle. This allows the model to naturally and robustly predict complex fracture phenomena, including the initiation of cracks from arbitrary sites, their subsequent propagation along complex paths, and events like crack branching and coalescence, all without the need for ad-hoc criteria or complex topological tracking and remeshing algorithms.
@@
The theoretical underpinnings of PF method are robust, originating from the variational theory of brittle fracture proposed by Francfort and Marigo, which recasts fracture as a global energy minimization problem Francfort and Marigo [1998]. The phase-field formulation can be rigorously shown to be a regularized approximation of this foundational theory Bourdin et al. [2000]. The great flexibility of the PF method framework, where the system’s behavior is governed by an energy functional, allows for its extension to coupled multi-physics problems by incorporating additional energy terms, see Quinteros et al. [2023], Zhuang et al. [2022], Asur Vijaya Kumar et al. [2022]. This capability has been leveraged to develop phase-field models for fracture in piezoelectric and ferroelectric materials, which, like the dielectric nanocomposites of interest here, exhibit strong electromechanical coupling Zhang et al. [2022], Quinteros et al. [2023]. These advanced models couple the variational formulation of fracture with the governing equations of electrostatics. A key challenge in this domain is the accurate modeling of the electrical boundary conditions on the crack faces, as the behavior depends on the electrical properties of the medium (e.g., air, vacuum) filling the crack gap. Different physical scenarios must be considered, such as an electrically impermeable (insulating) crack, a permeable crack, or a semi-permeable crack Ou and Chen [2003]. In the PF method framework, these conditions are encoded by selectively degrading different terms within the total electromechanical enthalpy density function based on the value of the phase-field variable Zhang et al. [2022], Quinteros et al. [2023]. A physically consistent model must therefore capture the simultaneous degradation of both the mechanical stiffness and the dielectric permittivity as the crack propagates, introducing significant nonlinearity and mathematical complexity into the governing equations.
@@
While the phase-field method provides the necessary physical fidelity to capture these complex, coupled phenomena, this power comes at a steep and often prohibitive price: extreme computational cost Cui et al. [2003]. This ”curse of fidelity” arises from several compounding factors. First, a standard PF method simulation already necessitates the use of a very fine computational mesh to accurately resolve the narrow, diffuse crack region, as the mesh size must be smaller than the phase-field length scale parameter. Second, when coupling with another physical field like electrostatics, a new set of governing equations and variables (e.g., electric potential) must be solved at every node in the mesh, increasing the size of the algebraic system. Third, and most critically, the coupling terms in the energy functional introduce strong nonlinearities, requiring more iterations within each time step for the numerical solver to converge. The combination of these factors leads to a multiplicative, rather than additive, increase in computational expense. As a result, high-fidelity, coupled electromechanical phase-field simulations can be so computationally intensive that they become impractical for tasks such as large-scale structural analysis, iterative design optimization, or real-time damage assessment, which are precisely the applications for which smart materials are intended. This establishes a critical research gap: a powerful theoretical tool exists to model the physics correctly, but it is too slow to be of practical use. There is a pressing need for a new computational paradigm that can retain the physical accuracy of coupled PF method while drastically reducing the associated computational burden.
@@
To surmount the computational bottleneck imposed by high-fidelity simulations, the scientific community is increasingly turning to machine learning (ML) and deep learning (DL). The central strategy is to develop a data-driven ”surrogate” model–a trained neural network that learns to approximate the complex -output relationship of the expensive physics-based simulation but can make predictions at a fraction of the computational cost Gao et al. [2025], Aldakheel et al. [2025], Liu et al. [2021], Herrmann and Kollmannsberger [2024a], Bahtiri et al. [2024], Dean et al. [2025]. The application of ML, particularly artificial neural networks (ANNs), to problems in fracture mechanics is a rapidly expanding field, see Gao et al. [2025], Aldakheel et al. [2025], Manav et al. [2024], Baktheer and Aldakheel [2025] among others. Early efforts have demonstrated the ability of ANNs to predict key fracture parameters, such as the stress intensity factor (SIF) for cracked plates, by training the networks on large datasets generated via conventional finite element analysis (FEA).
@@
For problems in continuum mechanics, where the outputs are spatially distributed fields (e.g., stress, strain, displacement, or phase-field), convolutional neural networks (CNNs) have emerged as a particularly suitable architecture. Originally developed for and famously successful in the domain of image recognition, CNNs are inherently designed to process grid-like data and learn hierarchical spatial features, see Tragoudas et al. [2024], Aldakheel et al. [2023], Herrmann and Kollmannsberger [2024b] and the references cited therein. The convolutional layers of a CNN act as learnable filters that slide across the input data, identifying local patterns. This mechanism is conceptually analogous to how physical phenomena, such as stress concentrations, manifest locally within a material domain, for instance, around the tip of a crack Shrivastava et al. [2022]. The efficacy of this approach has been demonstrated in several studies where CNNs have been successfully employed as surrogate models to predict the effective mechanical properties of periodic composites, see Messner [2019], Béji et al. [2023], Gavallas et al. [2024] among many others. In these works, the material microstructure is represented as a binary image, and the CNN is trained to regress the homogenized properties, achieving prediction speeds that are orders of magnitude faster than direct FEA simulations.
Recently, a growing body of research has focused on hybrid frameworks that combine phase-field fracture models with machine learning or deep learning techniques to accelerate fracture simulations or to infer damage evolution from high-fidelity data. In these studies, deep neural networks have been trained as surrogate solvers to approximate the solution of the phase-field governing equations, significantly reducing computational cost while retaining acceptable accuracy. Representative works include the use of convolutional neural networks to predict phase-field evolution in brittle fracture problems, recurrent or encoder–decoder architectures for time-dependent crack growth, and physics-informed or data-driven approaches to identify fracture parameters and crack paths in purely mechanical settings, see Gao et al. [2025], Aldakheel et al. [2025], Manav et al. [2024] among many others. These contributions have demonstrated the promise of PF–DL coupling for fast fracture prediction, particularly in elastic and quasi-brittle materials under mechanical loading.
In contrast to existing works, the present study advances the state of the art in two fundamental directions. First, we develop a hybrid deep learning framework built upon a fully coupled electromechanical phase-field formulation for dielectric nanocomposites, enabling the surrogate model to learn fracture behavior in a genuinely multi-physics setting. This allows the network to implicitly capture the simultaneous mechanical degradation and dielectric permittivity evolution induced by crack growth–a feature essential for modeling smart, self-sensing composite materials. Second, we introduce the electric potential field as a novel physical signature to guide the training of crack segmentation models. Rather than training the network solely on local strain or phase-field variables, we demonstrate that leveraging the electric potential–as a global and smoothly varying physical response–provides richer gradient information during the learning process. This leads to superior convergence, accuracy, and generalization, even when the final model performs predictions based solely on geometric configurations. To the authors’ knowledge, this is the first study to systematically exploit electric potential fields as the primary physical signal to inform geometry-to-crack mapping within a PF-DL framework.
@@
The accurate and efficient prediction of fracture in multifunctional dielectric nanocomposites is essential for their reliable design and deployment in next-generation technologies, particularly for real-time structural health monitoring. While coupled electromechanical phase-field models possess the requisite physical fidelity to capture the complex failure behavior of these materials, their prohibitive computational expense renders them impractical for the very applications they are meant to enable. This creates a critical disconnect between modeling capability and practical utility. To bridge this gap, this study proposes and validates a novel hybrid computational framework that synergistically combines high-fidelity, fully coupled electromechanical phase-field fracture modeling with deep learning-based surrogate modeling. By training a deep neural network on data generated from the physics-based PF method, the framework aims to achieve the predictive accuracy of the high-fidelity model at a computational speed amenable to large-scale analysis and design. The primary objectives and specific contributions of this paper are fourfold: (i) to develop and implement a thermodynamically consistent electromechanical phase-field fracture model for dielectric nanocomposites, capable of capturing the coupled degradation of mechanical stiffness and dielectric permittivity under electromechanical loading, (ii) to utilize this high-fidelity simulation framework to generate a comprehensive, high-resolution dataset of simulation results. This data set, which encompasses the evolution of phase-field and electric potential variables, serves as the ground-truth data for training the deep learning surrogate model, (iii) to design and train a deep CNN surrogate based on ResNet-U-Net architectures that learns to map initial geometric configurations (random defect patterns) directly to final crack paths. This stage leverages two distinct training strategies–one informed by phase-field data and the other by electric potential fields arising from electromechanical coupling–to assess which physical signature better guides the network in learning the underlying fracture mechanics, and (iv) to conduct a rigorous evaluation of the surrogate model’s performance, assessing its predictive accuracy on unseen scenarios and generalization capability.